# Analog In-memory Training on General Non-ideal Resistive Elements: The Impact of Response Functions

### v3

@@ I.4 Proof of Lemma 7: Descent of sequence $P_{k}$
$C_{k,+}:=$ $\frac{1}{2}\left(\max_{i\in[D]}\{q_{s}([P_{k}]_{i})\}+\min_{i\in[D]}\{q_{s}([P_{k}]_{i})\}\right),$ (111)
$C_{k,-}:=$ $\frac{1}{2}\left(\max_{i\in[D]}\{q_{s}([P_{k}]_{i})\}-\min_{i\in[D]}\{q_{s}([P_{k}]_{i})\}\right).$ (112)
@@
I.5 Proof of Corollary 1: Exact convergence of Tiki-Taka
@@
Optimizer. The digital SGD optimizer is implemented by FloatingPointRPUConfig in AIHWKIT, which is equivalent to the SGD implemented in PyTorch. The Analog SGD is implemented by selecting SingleRPUConfig as configuration and Tiki-Taka optimizers are implemented by UnitCellRPUConfig with TransferCompound devices in AIHWKIT.
Hardware. We conduct our experiments on one NVIDIA RTX 3090 GPU, which has 24GB memory and a maximum power of 350W. The simulations take from 30 minutes to 5 hours, depending on model sizes and datasets.
@@
The power response is a power function, given by
@@
We conduct training simulations of image classification tasks on a series of real datasets. In the implementation of Tiki-Taka, only a few columns or rows of $P_{k}$ are transferred per time to $W_{k}$ in the recursion (11) to balance the communication and computation. In our simulations, we transfer 1 column every time. The response granularity is set as $\Delta w_{\min}=$1e-3.
The other setting follows the settings of AIHWKIT, including output noise (0.5 % of the quantization bin width), quantization and clipping (output range set 20, output noise 0.1, and input and output quantization to 8 bits). Noise and bound management techniques are used in [71]. A learnable scaling factor is set after each analog layer, which is updated using SGD.
3-FC / MNIST. Following the setting in [16], we train a model with 3 fully connected layers. The hidden sizes are 256 and 128. The activation functions are Sigmoid. The learning rates are $\alpha=0.1$ for Digital SGD, $\alpha=0.05,\beta=0.01$ for Analog SGD and Tiki-Taka. The batch size is 10 for all algorithms. In Figure 4, the power response functions with $\gamma_{\text{res}}=0.5$ are used, and various $\tau$ are used as indicated in the legend.
CNN / MNIST. We train a convolution neural network, which contains 2-convolutional layers, 2-max-pooling layers, and 2-fully connected layers. The activation functions are Tanh. The first two convolutional layers use $\times5$ kernels with 16 and 32 kernels, respectively. Each convolutional layer is followed by a subsampling layer implemented by the max pooling function over non-overlapping pooling windows of size 2 $\times$ 2. The output of the second pooling layer, consisting of 512 neuron activations, feeds into a fully connected layer consisting of 128 tanh neurons, which is then connected into a 10-way softmax output layer. The learning rates are set as $\alpha=0.1$ for Digital SGD, $\alpha=0.05,\beta=0.01$ for Analog SGD are Residual Learning/Tiki-Taka. The batch size is 8 for all algorithms. In Figure 4, the power response functions with $\gamma_{\text{res}}=0.5$ are used, and various $\tau$ are used as indicated in the legend.
ResNet & MobileNet / CIFAR10 & CIFAR100. We train different models from the ResNet family, including ResNet18, 34, and 50. The base model is pre-trained on ImageNet dataset. The last fully connected layer is replaced by an analog layer. The learning rates are set as $\alpha=0.075$ for Digital SGD, $\alpha=0.075,\beta=0.01$ for Analog SGD, Residual Learning/Tiki-Taka, Tiki-Taka v2, and Residual Learningv2. Tiki-Taka adopts $\gamma=0.4$ unless stated otherwise. The batch size is 128 for all algorithms.
@@
CIFAR10/CIFAR100 ResNet. We fine-tune three models from the ResNet family with different scales on CIFAR10/CIFAR100 datasets. The power response functions with $\gamma_{\text{res}}=3.0$ and $\tau=0.1$, and the exponential response functions with $\gamma_{\text{res}}=4.0$ and $\tau=0.1$ are used, whose results are shown in Table 1 and 3, respectively. The results show that the Tiki-Taka outperforms Analog SGD by about $1.0\%$ in most of the cases in ResNet34/50, and the gap even reaches about $10.0\%$ for ResNet18 training on the CIFAR100 dataset.
@@
CIFAR10/CIFAR100 MobileNet. We fine-tune three MobileNet models with different scales on CIFAR10/CIFAR100 datasets. The response function is set as the power response with the parameter $\gamma_{\text{res}}=4.0$ and $\tau=0.05$, whose results are shown in Table 4. In the simulations, the accuracy of Analog SGD drops significantly by about $10\%$ in most cases, while Tiki-Taka remains comparable to the Digital SGD with only a slight drop.
@@
K.7 Ablation study on $\gamma$
We conduct a series of simulations to study the impact of mixing coefficient $\gamma$ in (10) on CIFAR10 or CIFAR100 dataset in the ResNet training tasks. The results are presented in Figure 6, which shows that Tiki-Taka achieves a great accuracy gain from increasing $\gamma$ from 0 to 0.1, while the gain saturates after that. Therefore, we conclude that Tiki-Taka benefits from a non-zero $\gamma$, and the performance is robust to the $\gamma$ selection.
Figure 6: The test accuracy of ResNet family models after 100 epochs trained by Tiki-Taka under different $\gamma$ in (10); (Left) CIFAR10. (Right) CIFAR100.

### v4

@@ I.4 Proof of Lemma 7: Descent of sequence $P_{k}$
$C_{k,+}:=$ $\frac{1}{2}\left(\max_{d\in[D]}\{q_{s}([P_{k}]_{d})\}+\min_{d\in[D]}\{q_{s}([P_{k}]_{d})\}\right),$ (110)
$C_{k,-}:=$ $\frac{1}{2}\left(\max_{d\in[D]}\{q_{s}([P_{k}]_{d})\}-\min_{d\in[D]}\{q_{s}([P_{k}]_{d})\}\right).$ (111)
@@
I.5 Proof of Corollary 1: Exact convergence of Residual Learning
@@
Optimizer. The baseline Digital SGD optimizer is implemented by FloatingPointRPUConfig in AIHWKIT, which is equivalent to the SGD implemented in PyTorch. The Analog SGD is implemented by selecting SingleRPUConfig as configuration, and Tiki-Taka optimizers are implemented by UnitCellRPUConfig with TransferCompound devices in AIHWKIT.
As suggested by [22], in the implementation of Residual Learning, only a few columns of $P_{k}$ are transferred per time to $W_{k}$ in the recursion (11) to balance the communication and computation. In our simulations, we transfer 1 column every time.
RPU Configuration. AIHWKIT offers fine-grained simulations of the hardware imperfections, such as the IO noise, analog-digital conversion, and so on. They are specified by the resistive processing unit (RPU) configurations. Without other specifications, we use the configuration list in Table˜4. The experimental setup uses a specific I/O configuration, as detailed in the relevant table. The system’s input and output signal bounds are explicitly defined. Regarding signal quality, the setup employs no input noise but introduces additive Gaussian noise to the output signal, the statistical properties of which are precisely specified. Finally, the resolution of the digital conversion processes is set by distinct bit values for both the input (DAC) and output (ADC).
In addition, noise, bound, and update management techniques are used in [71]. A learnable scaling factor is applied after each analog layer and updated using SGD. For each gradient update step, if more than $\text{BL}=32$ pulses are desired, only BL pulses are fired.
Table 4: Hardware imperfection setting
configuration value
input bound $1.0$
input noise None
input resolution (DAC) 77 bits
output bound $12.0$
output noise additive Gaussian noise $\mathcal{N}(0,0.06^{2})$
output resolution (ADC) 99 bits
Update granularity $\Delta w_{\min}$ $1\times 10^{-3}$
Bit length BL $32$
Simulation hardware. We conduct our experiments on one NVIDIA RTX 3090 GPU, which has 24GB of memory and a maximum power of 350W. The simulations take from 30 minutes to 5 hours, depending on model sizes and datasets.
@@
We consider two types of response functions in our simulations: power and exponential response functions with dynamic ranges $\tau,\tau]$, The power response is a power function, given by
@@
We conduct training simulations of image classification tasks on a series of real datasets.
3-FC @ MNIST. Following the setting in [16], we train a model with 3 fully connected layers. The hidden sizes are 256 and 128. The activation functions are Sigmoid. The learning rates are $\alpha=0.1$ for Digital SGD, $\alpha=0.05,\beta=0.01$ for Analog SGD and Tiki-Taka. The batch size is 10 for all algorithms. In Figure 4, the power response functions with $\gamma_{\text{res}}=0.5$ are used, and various $\tau$ are used as indicated in the legend.
CNN @ MNIST. We train a convolutional neural network, which contains 2 convolutional layers, 2 max-pooling layers, and 2 fully connected layers. The activation functions are Tanh. The first two convolutional layers use $\times5$ kernels with 16 and 32 kernels, respectively. Each convolutional layer is followed by a subsampling layer implemented by the max pooling function over non-overlapping pooling windows of size 2 $\times$ 2. The output of the second pooling layer, consisting of 512 neuron activations, feeds into a fully connected layer consisting of 128 tanh neurons, which is then connected to a 10-way softmax output layer. The learning rates are set as $\alpha=0.1$ for Digital SGD, $\alpha=0.05,\beta=0.01$ for Analog SGD are Residual Learning/Tiki-Taka. The batch size is 8 for all algorithms. In Figure 4, the power response functions with $\gamma_{\text{res}}=0.5$ are used, and various $\tau$ are used as indicated in the legend.
ResNet/MobileNet @ CIFAR10/CIFAR100. We train different models from the ResNet family, including ResNet18, 34, and 50. The base model is pre-trained on the ImageNet dataset. The last fully connected layer is replaced by an analog layer. The learning rates are set as $\alpha=0.075$ for Digital SGD, $\alpha=0.075,\beta=0.01$ for Analog SGD, Residual Learning/Tiki-Taka, Tiki-Taka v2, and Residual Learningv2. Tiki-Taka adopts $\gamma=0.4$ unless stated otherwise. The batch size is 128 for all algorithms.
@@
ResNet @ CIFAR10/CIFAR100. We fine-tune three models from the ResNet family with different scales on CIFAR10/CIFAR100 datasets. The power response functions with $\gamma_{\text{res}}=3.0$ and $\tau=0.1$, and the exponential response functions with $\gamma_{\text{res}}=4.0$ and $\tau=0.1$ are used, whose results are shown in Table 1 and 5, respectively. The results show that the Tiki-Taka outperforms Analog SGD by about $1.0\%$ in most of the cases in ResNet34/50, and the gap even reaches about $10.0\%$ for ResNet18 training on the CIFAR100 dataset.
@@
MobileNet @ CIFAR10/CIFAR100. We fine-tune three MobileNet models with different scales on CIFAR10/CIFAR100 datasets. The response function is set as the power response with the parameter $\gamma_{\text{res}}=4.0$ and $\tau=0.05$, whose results are shown in Table 6. In the simulations, the accuracy of Analog SGD drops significantly by about $10\%$ in most cases, while Tiki-Taka remains comparable to the Digital SGD with only a slight drop.